# Benchmarking Retrieval-Augmented Large Language Models in Biomedical NLP: Application, Robustness, and Self-Awareness

### v2

@@
Abstract
Large language models (LLM) have demonstrated remarkable capabilities in various biomedical natural language processing (NLP) tasks, leveraging the demonstration within the input context to adapt to new tasks. However, LLM is sensitive to the selection of demonstrations. To address the hallucination issue inherent in LLM, retrieval-augmented LLM (RAL) offers a solution by retrieving pertinent information from an established database. Nonetheless, existing research work lacks rigorous evaluation of the impact of retrieval-augmented large language models on different biomedical NLP tasks. This deficiency makes it challenging to ascertain the capabilities of RAL within the biomedical domain. Moreover, the outputs from RAL are affected by retrieving the unlabeled, counterfactual, or diverse knowledge that is not well studied in the biomedical domain. However, such knowledge is common in the real world. Finally, exploring the self-awareness ability is also crucial for the RAL system. So, in this paper, we systematically investigate the impact of RALs on 5 different biomedical tasks (triple extraction, link prediction, classification, question answering, and natural language inference). We analyze the performance of RALs in four fundamental abilities, including unlabeled robustness, counterfactual robustness, diverse robustness, and negative awareness. To this end, we proposed an evaluation framework to assess the RALs’ performance on different biomedical NLP tasks and establish four different testbeds based on the aforementioned fundamental abilities. Then, we evaluate 3 representative LLMs with 3 different retrievers on 5 tasks over 9 datasets. The evaluation indicates that while RALs enhance the performance of most biomedical datasets used and demonstrate a degree of counterfactual robustness, they still encounter significant challenges with unlabeled and counterfactual retrieval information, as well as negative awareness.
Lately, significant progress has been made in large language models (LLMs) such as ChatGPT  [^1] https://chat.openai.com/. To adapt the LLM to the biomedical domain, several LLMs have been developed, such as MedLLaMA-13B Wu et al. (2023) and Med-PaLM 2 Singhal et al. (2023). Despite demonstrating impressive general capabilities Li & Zhang (2023), these models still face significant challenges, including factual hallucination  Ji et al. (2023); Zhang et al. (2023) , and absence of newly uploaded knowledge Ovadia et al. (2023).
@@ Abstract
Retrieval-augmented language models Li & Zhang (2023); Lewis et al. (2020); Li & Huang (2023), in contrast, can retrieve knowledge from an external datastore when needed, potentially reducing hallucination and improve the new knowledge adaption ability. The most common method is to use the designed retriever to retrieve the knowledge that is relevant to the input sentence, subsequently, the retrieved knowledge, along with the input sentence, is fed into the LLM to assist in generating the expected output.
In the question answering (QA) task, a retrieval-augmented language model can access knowledge from unlabeled corpus [^2] In this work, corpus refers to the knowledge base that needs to be retrieved. like PubMed. The QA format allows the unlabeled corpus to potentially furnish answers to questions. However, for tasks like triple extraction, incorporating the unlabeled corpus may yield adverse effects. Counterfactual information, such as error annotations, is prevalent in labeled corpora, presenting challenges for retrievers in obtaining useful information. Additionally, LLMs still grapple with generating unreliable information retrieved incorrectly. Incorporating diverse knowledge holds promise for improving model performance. For example, question answering relies on extracting information from extensive contexts, thus potentially impacting information extraction performance. Moreover, the influence of retrieval information from various tasks or datasets on RAL performance remains underexplored. Self-awareness is crucial for RALs, if RALs can distinguish between positive retrieval knowledge and negative knowledge, they have the opportunity to rectify their actions.
These challenges hinder RALs from consistently producing reliable and accurate responses. Unfortunately, in the biomedical domain, only a few studies, such as Almanac Zakka et al. (2024), have explored the RAL performance in QA, leaving a gap in understanding how these factors affect RAL across various biomedical NLP tasks. Consequently, there is a pressing need for a comprehensive evaluation of RALs with different LLMs across biomedical NLP tasks. To this end, this paper conducts a comprehensive evaluation of RAL for different LLMs on 5 biomedical NLP tasks over 9 datasets. Specifically, we create a new RAL benchmark for biomedical NLP, namely BioRAB, as shown in Figure 1, and create 4 testbeds to evaluate the mentioned fundamental abilities.
@@
Utilizing BioRAB, we evaluate its performance across 5 tasks (triple extraction, link prediction, text classification, question answering, and natural language inference) using 9 biomedical NLP datasets. Furthermore, BioRAB undergoes evaluation with three widely used LLMs: LLaMA2-13B, MedLLaMA-13B, and LLaMA3-8B, utilizing three commonly employed retrievers (BM25, Contriever, and MedCPT).
We observed that although RALs can enhance response accuracy in the majority of biomedical NLP tasks we evaluated, they encounter notable challenges. Particularly in the question-answering task, we noted that RALs did not yield significant improvements in the datasets we used. We speculate that this could be attributed to the limitations of the corpus used for retrieving, as the training dataset (corpus we used for retrieving) may not have provided sufficient information compared to using Wikipedia or PubMed. Moreover, RALs struggle to generate the desired output when the corpus lacks labeling when compared to the labeled corpus. An interesting finding is that in datasets like ChemProt and Hetionet, RALs exhibit improved performance with unlabeled corpora compared to the source LLM. Besides, RALs lack the capability to extract useful information from counterfactual corpora and struggle to discern the most relevant information. We also find this is not a common case, some datasets, such as in the dataset ADE and Hetionet, RAL could handle the counterfactual instance. Additionally, when presented with a diverse labeled corpus, RALs do not achieve optimal performance across tasks, except for the natural language inference task. Finally, we found despite providing counterfactual examples during training, the LLM was still able to generate correct outputs in some instances. However, RALs struggle with self-awareness, as they lack the ability to determine which examples could help improve model performance. The experimental results mentioned above underscore the necessity for further resolution of these issues for the RAL. Our contributions are the following:
@@
  • We evaluate the RAL on 5 different biomedical tasks over 9 datasets by using 3 LLMs with 3 retrievers.
1 Related Work
1.1 Retrieval-Augmented Language Models (RALMs)
Many studies Li & Zhang (2023); Lewis et al. (2020); Guu et al. (2020); Ram et al. (2023); Li et al. (2024a; b), have been proposed to use retrieved information from various knowledge stores to better understand the text or generate the expected output. For example, KIEST Li & Huang (2023) dynamically injects retrieved entity and attribute knowledge from the knowledge graph when generating the entity or attribute in the task of entity stage changes. Lewis et al. (2020) uses the maximum inner product search (MIPS) to find the top-K documents which are combined with a query to predict the answers. To enhance retrieval capability, BiomedRAG Li et al. (2024a) proposes a learned retriever to retrieve the chunk information from the build database and improve the model performance. while CTLP Li et al. (2024b) aims to create a condensed transition graph to improve the link prediction performance, the sample paths between two entities are retrieved to construct the condensed transition graph. RT Li & Zhang (2023) employs a chain of thought and retrieves pertinent labeled sentences to enhance few-shot biomedical named entity recognition (NER) tasks. However, there is no work to systematically evaluate the effectiveness of retrieval-argument LLM on the different biomedical NLP tasks.
1.2 Evaluation of RAL
Evaluating RALs has received significant attention due to their remarkable general capability. It enables researchers to gain a deeper understanding of the limitations and abilities of LLMs. However, there are few studies Zakka et al. (2024); Xiong et al. (2024) focusing on the evaluation of RAGs in the biomedical domain, primarily centered around question answering tasks. For example, Xiong et al. Xiong et al. (2024) evaluated RAG models on five biomedical QA datasets using the zero-shot setting. Almanac Zakka et al. (2024) evaluates the ChatGPT with one retriever and one QA dataset. In contrast to the current work, we offer a broader evaluation across four testbeds and 5 tasks spanning 9 datasets.
2 BioRAB: Biomedical Retrieval-Augmented Generation Benchmark
In this section, we begin by outlining the operational flow of the RALs. Following this, we introduce the proposed four abilities and the building progress of four relevant testbeds. Finally, we introduce the evaluation metrics employed to assess performance.
2.1 RAL Working Flow
To solve the hallucination problem, RAL is proposed to retrieve the external knowledge from the corpus and improve the LLM performance. Generally, as shown in Figure 2(a), retrieved corpus needs to be constructed initially, in numerous question-answering RAL models, the corpus primarily originates from the unlabeled open source such as PubMed, Textbook. However, for some label-sensitive tasks, such as triple extraction, the unlabeled open source may be invalid. In our work, the corpus is defined as the training set for the relevant task. For instance, as illustrated in Figure 2(a), if "n" denotes the PHarmKG, each key corresponds to a sentence what is the relationship between the head entity and tail entity? in its training set, while the corresponding value denotes the relevant label relationship for that key. In the second step, the retriever is used to obtain the relevant (key, value) pairs from the corpus based on the input sentence. At last, the retrieved (key, value) pairs with the input sentence are fed into the LLM to generate the expected output. For each instance $X$ of each "n", there are three components: Instruction $I$, context $C$, and response $R$. For example, in the training dataset of ade-corpus-v2 (classification task), if the label of a sentence $S$: She had been administered tacrolimus for prophylaxis of graft-versus-host reaction is False in the $X$, $I=$You are an excellent linguist. The task is to predict whether this sentence is True or False. Examples: context: The hemangioma regressed markedly 6 weeks after the procedure and serous retinal detachment showed marked resolution.response: False, $C=S$, $R=False$.
2.2 Defined four Ability of BioRAL
2.2.1 unlabeled Robustness (UR)
Not all tasks have vast labeled retrieval corpus, therefore, for each task, the retriever must gather information from unlabeled corpora, while the RAL may still have the ability to generate the expected results. To evaluate the efficacy of RAL in this regard, we introduce our proposed UR testbed. Specifically, as shown in Figure 2(b), the corpus of "n" is defined as the training set without value(label) for the "n". The retriever retrieves the relevant information from this unlabeled corpus. After that, the retrieved Key with the input sentence is fed into the LLM. For example, in the training dataset of ade-corpus-v2 (classification task), if the label of a sentence $S$: She had been administered tacrolimus for prophylaxis of graft-versus-host reaction is False In the $X$, $I=$You are an excellent linguist. The task is to predict whether this sentence is True or False, retrieved sentence: A macrophage activation syndrome, possibly related to methotrexate toxicity, developed in a boy with systemic juvenile rheumatoid arthritis, $C=S$, $R=False$.
2.2.2 Counterfactual Robustness (CR)
2.2.3 Diverse Robustness (DR)
2.2.4 Negative Awareness (NA)
2.3 Evaluation Metrics
2.3.1 Task-based Metrics
In the triple Extraction task, same as BiomedRAG Li et al. (2024a), triple is regarded as correct when its relation type, the head entity, and the tail entity are all correct. For example, in the sentence: Infusion of prostacyclin (PGI2) reportedly attenuates renal ischemic injury in the dog and the rat., triple <Infusion, treats, rat> is regarded as correct while <injury, treats, rat> is not. We evaluated all the models and reported the evaluation metric, including Micro Precision, Recall, and F1-score. For the text classification, link prediction, and question answering task, we follow the same evaluation metrics as triple extraction. For the natural language inference task, we use the same evaluation metric (Macro F1) as the BioNLI.
2.3.2 Negative Awareness Metrics
3 Experiments
In this section, we assess RAL’s performance across various biomedical NLP tasks, analyze its efficacy on four proposed testbeds, and discuss its abilities.
3.1 Settings and Dataset
We evaluated three state-of-the-art LLMs: LLamA2-13B (Touvron et al., 2023) , MedLLamA-13B (Wu et al., 2023), and LLaMA3 8B [^3] https://github.com/meta-llama/llama3, along with three retrievers: BM25 (Luo et al., 2023), Contriver (Izacard et al., 2021), and MedCPT (Jin et al., 2023). We considered five biomedical NLP tasks: triple extraction, link prediction, text classification, question answering, and natural language inference, across nine datasets: ADE, ChemProt, GIT, PHarmKG, Hetionet, Ade-corpus-v2, SemedCLass, MedMCQA, and BioNLI. The data statistics are shown in Table 1. The experiments were conducted using A100 GPUs.
Dataset train test dev
Triple extraction ADE (Gurulingappa et al., 2012b) 4,970 2,130
ChemProt (Taboureau et al., 2010) 4,001 3,355 2,366
GIT (Li et al., 2023) 3,734 465 492
Link Prediction PHarmKG (Zheng et al., 2021) 4,000 500 500
Hetionet (Himmelstein et al., 2017) 4,000 500 500
Text classification Ade-corpus-v2 (Gurulingappa et al., 2012a) 4,000 500 500
SemdClass (Vasilakes Jake A, 2018) 2,400 600 600
Question answering MedMCQA (Pal et al., 2022) 34,994 4,183 4,183
Natual language inference BioNLI (Bastan et al., 2022) 5,544 6,308 12,807
Table 1: Data Statistics for the datasets we used in this work
3.1.1 Triple Extraction Dataset
  • ADE (Gurulingappa et al., 2012a) is extended from relation extraction task to triplet extraction task in this paper. All sentences either describe the effect of the drug or the dose of the drug. Thus, the triplets consist of (head entity: drug, relation type: effect, tail entity: effect_description) and (head entity: drug, relation type: dosage, tail entity: dose_description). Among all triplets, there are only two relation types: effect and dosage.
  • ChemProt (Taboureau et al., 2010): The Chemical Protein Interaction Corpus comprises 2432 PubMed abstracts annotated with chemical-protein interactions, encompassing 23 distinct interaction relations. Building upon prior research (Sun et al., 2022), the corpus exclusively considers sentence-level instances, with a particular focus on five prominent interaction types for classification: CPR3, CPR4, CPR5, CPR6, CPR9.
  • GIT (Li et al., 2023) is a high-quality biomedical triple extraction dataset for non-drug therapies, characterized by its high-quality annotations and comprehensive coverage of relation types. It includes 22 relation types from SemMedDB.
3.1.2 Link Prediction
  • PHarmKG (Zheng et al., 2021) is a knowledge graph to describe the relationship among genes, drugs, and diseases. In this work, we aim to predict the four mentioned relation types (Interactions, Disease-Gene, Disease-Chemical, Chemical-Gene) between two entities. During the huge quantity of triples in the PHarmKG, we randomly select 4,000 samples from the source training set for training, 500 samples from the source testing set for testing, and 500 samples from the source validation set for validation.
  • Hetionet (Himmelstein et al., 2017) is an integrative network of disease, which includes 46 relation types. In our paper, we randomly select 4,000 samples from the source training set for training, 500 samples from the source testing set for testing, and 500 samples from the source validation set for validation.
3.1.3 Text Classification
  • Ade-corpus-v2 (Gurulingappa et al., 2012a) dataset is designed for classifying whether a sentence is ADE( Adverse Drug Reaction)-related (True) or not (False). In our paper, we randomly select 4,000 instances for training, 500 for testing, and 500 for validation.
  • The SemdClass (Vasilakes Jake A, 2018), aims to understand whether the provided triple belongs to the given sentence or not. It includes two classes, False and True.
3.1.4 Questing Answering and Natual Language Inference
  • MedMCQA (Pal et al., 2022) is a multi-choice question-answering dataset that designed to address the medical entrance exam questions. In this work, we opt for the five-choice version (A, B, C, D, E).
  • BioNLI (Bastan et al., 2022) aims to understand whether the provided hypothesis is consistent or adversarial to the premise.
3.2 Comparison between RALs with backbone LLMs
@@
We first benchmark various LLMs and RALs on 9 datasets, the results are shown in Table 2 and Table 3. In the triple extraction task, we observed that RALs outperformed LLMs (specifically RALs without a retriever), achieving better performance. For example, RALs (MedLLaMA 13B with Contriever) enhanced the original MedLLaMA 13B by 22.37%, in terms of F1 score on the ADE dataset. However, RAL still faces challenges in entity recognition. For example, in ADE, LLaMA2-13B gets the best performance when compared to LLaMA2-13B with retrievers. Another interesting finding is that models with bigger parameters may not necessarily yield the best performance. For instance, on the ChemProt, LLaMA3-8B with Contriver outperforms other RALs with larger parameter sizes.
@@
On MedMCQA, our findings differ from other works Xiong et al. (2024) as we observed that LLMs outperform RALs in achieving the best performance, we speculate that the reason for this discrepancy lies in the nature of label-not-sensitive tasks, where RALs have the capability to retrieve large corpora such as PubMed White (2020) or other relevant datasets. In our study, however, our corpus is derived solely from the training set, which may limit the breadth of knowledge accessible to the RALs.
triple head entity relation tail entity
ADE LLaMA2-13B BM25 (Luo et al., 2023) 30.99 30.88 30.93 73.96 73.71 73.84 94.85 94.54 94.70 49.10 48.94 49.02
Contriever (Izacard et al., 2021) 36.07 36.06 36.06 79.76 79.72 79.74 93.80 93.76 93.78 48.99 48.97 48.98
MedCPT (Jin et al., 2023) 30.81 30.80 30.81 76.75 76.71 76.73 94.41 94.37 94.39 43.87 43.85 43.86
MedLLaMA-13B BM25 (Luo et al., 2023) 33.77 33.76 33.77 77.06 77.02 77.04 94.82 94.77 94.80 51.20 51.18 51.19
Contriever (Izacard et al., 2021) 35.66 33.57 34.58 79.15 74.51 76.76 94.36 88.83 91.51 49.18 46.29 47.69
MedCPT (Jin et al., 2023) 33.30 29.72 31.41 79.01 70.52 74.52 95.48 85.21 90.05 45.77 40.85 43.17
LLaMA3-8B BM25 (Luo et al., 2023) 27.88 27.70 27.79 72.87 72.39 72.63 94.61 93.99 94.30 45.79 45.49 45.64
Contriever (Izacard et al., 2021) 34.44 31.13 32.70 78.91 71.31 74.92 93.14 84.18 88.43 48.57 43.90 46.12
MedCPT (Jin et al., 2023) 31.70 27.04 29.19 77.99 66.53 71.80 94.00 80.19 86.55 45.35 38.69 41.75
ChemProt LLaMA2-13B BM25 (Luo et al., 2023) 49.44 48.78 49.11 65.77 64.89 65.33 75.56 74.55 75.05 65.51 64.63 65.07
Contriever (Izacard et al., 2021) 85.75 85.05 85.40 98.42 97.61 98.01 91.58 90.84 91.21 94.02 93.25 93.63
MedCPT (Jin et al., 2023) 86.25 85.40 85.82 98.44 97.47 97.95 91.19 90.29 90.73 95.53 94.59 95.06
MedLLaMA-13B BM25 (Luo et al., 2023) 54.78 49.53 52.02 73.30 66.29 69.62 78.10 70.62 74.17 74.40 67.28 70.66
Contriever (Izacard et al., 2021) 86.15 85.22 85.69 97.97 96.92 97.44 91.38 90.40 90.89 95.18 94.15 94.66
MedCPT (Jin et al., 2023) 81.33 80.08 80.70 98.14 96.63 97.38 89.60 88.22 88.91 91.96 90.55 91.25
LLaMA3-8B BM25 (Luo et al., 2023) 70.23 69.98 70.10 83.86 83.57 83.71 84.30 84.00 84.15 91.54 91.22 91.38
Contriever (Izacard et al., 2021) 87.13 86.68 86.91 97.92 97.41 97.67 91.26 90.78 91.02 95.56 95.06 95.30
MedCPT (Jin et al., 2023) 86.62 83.07 84.81 98.06 94.04 96.01 92.54 88.74 90.60 94.06 90.20 92.09
GIT LLaMA2-13B BM25 (Luo et al., 2023) 60.81 54.73 57.61 76.94 69.25 72.89 76.70 69.03 72.67 77.78 70.00 73.68
Contriever (Izacard et al., 2021) 74.78 72.37 73.55 89.44 86.56 87.98 83.22 80.54 81.86 87.11 84.30 85.68
MedCPT (Jin et al., 2023) 75.64 73.44 74.52 92.03 89.35 90.67 83.06 80.65 81.83 89.04 86.45 87.73
MedLLaMA-13B BM25 (Luo et al., 2023) 58.59 57.20 57.89 79.30 77.42 78.35 72.91 71.18 72.03 78.63 76.77 77.69
Contriever (Izacard et al., 2021) 65.95 65.81 65.88 85.34 85.16 85.25 75.86 75.70 75.78 83.84 83.66 83.75
MedCPT (Jin et al., 2023) 75.65 74.84 75.24 90.22 89.25 89.73 82.39 81.51 81.95 90.00 89.03 89.51
LLaMA3-8B BM25 (Luo et al., 2023) 62.72 62.58 62.65 79.31 79.14 79.22 79.74 79.57 79.66 77.80 77.63 77.72
Contriever (Izacard et al., 2021) 74.76 67.53 70.96 89.76 81.08 85.20 82.14 74.19 77.97 88.33 79.78 83.84
MedCPT (Jin et al., 2023) 66.20 50.97 57.59 87.15 67.10 75.82 82.96 63.87 72.17 78.77 60.65 68.53
Table 2: Results of various approaches for triple extraction on ADE, ChemProt, and GIT. Underline with shade (green, pink, and blue) indicates the best performance on ADE, ChemProt, and GIT separately.
@@
PHarmKG Hetionet Ade-corpus-v2 SemClass MedMCQA BioNLI
LLaMA2-13B BM25 (Luo et al., 2023) 97.60 97.60 97.60 82.37 82.37 82.37 95.40 95.40 95.40 75.50 75.50 75.50 40.38 40.49 40.42 45.10
Contriever (Izacard et al., 2021) 98.00 98.00 98.00 77.00 77.00 77.00 96.60 96.60 96.60 79.33 79.33 79.33 35.53 35.52 35.52 35.12
MedCPT (Jin et al., 2023) 97.40 97.40 97.40 81.60 81.60 81.60 96.80 96.80 96.80 78.50 78.50 78.50 36.78 36.93 36.80 69.21
MedLLaMA-13B BM25 (Luo et al., 2023) 95.00 95.00 95.00 90.40 90.40 90.40 95.60 95.60 95.60 72.67 72.67 72.67 37.81 37.96 37.86 48.81
Contriever (Izacard et al., 2021) 97.00 97.00 97.00 77.20 77.20 77.20 95.60 95.60 95.60 77.66 77.66 77.66 29.82 29.75 29.77 53.07
MedCPT (Jin et al., 2023) 97.40 97.40 97.40 84.40 84.40 84.40 95.40 95.40 95.40 76.16 76.16 76.16 33.86 34.04 33.88 53.68
LLaMA3-8B BM25 (Luo et al., 2023) 96.80 96.80 96.80 81.80 81.80 81.80 94.80 94.80 94.80 75.50 75.50 75.50 37.73 38.90 37.79 19.17
Contriever (Izacard et al., 2021) 96.60 96.60 96.60 73.40 73.40 73.40 94.60 94.60 94.60 75.83 75.83 75.83 28.11 28.12 28.11 63.85
MedCPT (Jin et al., 2023) 97.00 97.00 97.00 83.00 83.00 83.00 95.40 95.40 95.40 74.67 74.67 74.67 31.57 31.82 31.56 56.89

### v3

@@
Abstract: To reduce hallucinations in large language models (LLMs), retrieval-augmented LLMs (RALs) retrieve supporting knowledge from external databases. However, their performance on biomedical natural language processing (NLP) tasks remains underexplored. We introduce BioRAB, a comprehensive evaluation framework assessing RALs across five biomedical NLP tasks and 11 datasets, using four testbeds: unlabeled robustness, counterfactual robustness, diverse robustness, and self-awareness. To improve RALs’ robustness and negative awareness, we propose a Detect-and-Correct strategy and a contrastive learning approach. Experimental results show that RALs generally outperform standard LLMs on most biomedical tasks, but still struggle with robustness and self-awareness, particularly under counterfactual and diverse scenarios. Our proposed methods significantly improve performance in robustness to unlabeled and counterfactual data, and increase the model’s ability to detect and avoid incorrect predictions. These findings highlight key limitations in current RALs and underscore the need for continued refinement to ensure reliability and accuracy in high-stakes biomedical applications.
Teaser: BioRAB evaluates retrieval-augmented language models on robustness and self-awareness in biomedical NLP tasks.
1 Introduction
Recently, significant progress has been made in large language models (LLMs). To adapt the LLM to the biomedical domain, several biomedical focused LLMs have been developed, such as MedLLaMA-13B [1] and Med-PaLM 2 [3]. Despite demonstrating impressive general capabilities [4], these models still face significant challenges, including factual hallucination  [5] , and absence of newly uploaded knowledge [6].
Retrieval-augmented language models (RALs) [4, 7, 8, 9], in contrast, can retrieve knowledge from an external datastore when needed, potentially reducing hallucination and improve the new knowledge adaption ability. The most common method is to use the designed retriever to retrieve the knowledge that is relevant to the input sentence, subsequently, the retrieved knowledge, along with the input sentence, is fed into the LLM to assist in generating the expected output.
In the question answering (QA) task, a retrieval-augmented language model can access knowledge from an unlabeled corpus, which serves as a retrievable knowledge base, such as PubMed. The QA format allows the unlabeled corpus to potentially furnish answers to questions. However, for tasks like triple extraction, incorporating the unlabeled corpus may yield adverse effects. Counterfactual information, such as error annotations, is prevalent in labeled corpora, presenting challenges for retrievers in obtaining useful information. Additionally, LLMs still grapple with generating unreliable information retrieved incorrectly. Incorporating diverse knowledge holds promise for improving model performance. For example, question answering relies on extracting information from extensive contexts, thus potentially impacting information extraction performance. Moreover, the influence of retrieval information from various tasks or datasets on RAL performance remains underexplored. Self-awareness is crucial for RALs, if RALs can distinguish between positive retrieval knowledge and negative knowledge, they have the opportunity to rectify their actions.
@@ 1 Introduction
These challenges hinder RALs from consistently producing reliable and accurate responses. Unfortunately, in the biomedical domain, only a few studies, such as Almanac [10], have explored the RAL performance in QA, leaving a gap in understanding how these factors affect RAL across various biomedical NLP tasks. Consequently, there is a pressing need for a comprehensive evaluation of RALs with different LLMs across biomedical NLP tasks. To this end, this paper conducts a comprehensive evaluation of RAL for different LLMs on 5 biomedical NLP tasks over 11 datasets. Specifically, we create a new RAL benchmark for biomedical NLP, namely BioRAB, as shown in Figure 1, and create 4 testbeds to evaluate the mentioned fundamental abilities.
  1. 1.
@@
Utilizing BioRAB, we evaluate its performance across 5 tasks (triple extraction, link prediction, text classification, question answering, and natural language inference) using 11 biomedical NLP datasets. Furthermore, BioRAB undergoes evaluation with five widely used LLMs: LLaMA2-13B [13] , MedLLaMA-13B [1], LLaMA3-8B [14], Phi4 14B [15] and Qwen2.5 32B [17], utilizing three commonly employed retrievers (BM25 [2], Contriever [11], and MedCPT [12].
We observed that although RALs can enhance response accuracy in the majority of biomedical NLP tasks we evaluated, they encounter notable challenges. Particularly in the question-answering task, we noted that RALs did not yield significant improvements in the datasets we used. We speculate that this could be attributed to the limitations of the corpus used for retrieving, as the training dataset (corpus we used for retrieving) may not have provided sufficient information compared to using Wikipedia or PubMed. Moreover, RALs struggle to generate the desired output when the corpus lacks labeling when compared to the labeled corpus. An interesting finding is that in datasets like ChemProt and Hetionet, RALs exhibit improved performance with unlabeled corpora compared to the source LLM. Besides, RALs lack the capability to extract useful information from counterfactual corpora and struggle to discern the most relevant information. We also find this is not a common case, some datasets, such as in the dataset ADE and Hetionet, RAL could handle the counterfactual instance. Additionally, when presented with a diverse labeled corpus, RALs do not achieve optimal performance across tasks, except for the natural language inference task. Finally, we found despite providing counterfactual examples during training, the LLM was still able to generate correct outputs in some instances. However, RALs struggle with self-awareness, as they lack the ability to determine which examples could help improve model performance.
To address these challenges, we propose a Detect-and-Correct method aimed at enhancing RAL’s ability to identify and rectify errors in the retrieved corpus, thereby improving its overall reliability. Additionally, we introduce a contrastive learning approach that refines RAL’s negative awareness by strengthening its capacity to distinguish differences of samples. By integrating these two techniques, our method significantly enhances RAL’s robustness and awareness ablilty, making it more effective in processing complex retrieval tasks. Together, these improvements contribute to a more resilient and accurate model for biomedical applications.
Our contributions are the following:
  1. 1.
@@
  2. 2.
    we propose a Detect-and-Correct method and a contrastive learning approach to enhance the RAL’s robustness (unlabeled and counterfactual) and negative awareness;
  3. 3.
@@
    We evaluate the RAL on 5 different biomedical tasks over 11 datasets by using 5 LLMs with 3 retrievers.
2 Results
2.1 Results of RALs and backbone LLMs
@@
We first benchmark various LLMs and RALs on 11 datasets, the results are shown in Table 1. In the triple extraction task, we observed that RALs outperformed LLMs (specifically RALs without a retriever) on LLaMA2-13B,MedLLaMA-13B, LLaMA3-8B and phi4 14B achieving better performance. For example, RALs (MedLLaMA 13B with Contriever) enhanced the original MedLLaMA 13B by 22.37%, in terms of F1 score on the ADE dataset. Another finding is that for Qwen2.5-32B, using different retrievers does not lead to significant performance improvement.
@@
On MedMCQA, our findings differ from other works [18] as we observed that LLMs outperform RALs in achieving the best performance, we speculate that the reason for this discrepancy lies in the nature of label-not-sensitive tasks, where RALs have the capability to retrieve large corpora such as PubMed [19] or other relevant datasets. In our study, however, our corpus is derived solely from the training set, which may limit the breadth of knowledge accessible to the RALs.
Triple Extraction Link Prediction Classification QA NL Inference
@@
LLM Approach ADE ChemProt GIT PHarmKG Hetionet DS Ade-corpus-v2 SemClass SDoH classification MedMCQA BioNLI
LLaMA2-13B BM25 [2] 30.93 49.11 57.61 97.60 82.37 75.86 95.40 75.50 63.71 40.42 45.10
Contriever [11] 36.06 85.40 73.55 98.00 77.00 76.50 96.60 79.33 61.18 35.52 35.12
MedCPT [12] 30.81 85.82 74.52 97.40 81.60 78.01 96.80 78.33 68.77 36.80 69.21
No Retriever 34.86 77.48 58.99 97.60 80.80 77.15 96.40 77.66 72.57 41.52 62.62
MedLLaMA-13B BM25 [2] 33.77 52.02 57.89 95.00 90.04 71.33 95.60 72.67 65.40 37.86 48.81
Contriever [11] 34.58 85.69 65.88 97.00 77.20 75.64 95.60 77.66 56.96 29.77 53.07
MedCPT [12] 31.41 80.70 75.24 97.40 84.40 72.84 95.40 76.16 66.66 33.88 53.68
No Retriever 12.21 50.52 42.05 97.20 78.54 80.38 95.40 64.00 67.71 46.47 61.07
LLaMA3-8B BM25 [2] 27.79 70.10 62.65 96.80 81.80 75.43 94.80 75.50 71.30 37.79 19.17
Contriever [11] 32.70 86.91 70.96 96.60 73.40 78.01 94.60 75.83 72.57 28.11 63.85
MedCPT [12] 29.19 84.81 57.59 97.00 83.00 76.29 95.40 74.67 71.72 31.56 56.89
No Retriever 7.05 21.32 74.27 97.20 81.80 80.38 93.80 73.16 57.00 55.91 6.71
Phi4 14B BM25 [2] 36.23 81.26 60.61 97.20 88.80 60.75 88.20 74.67 78.90 35.49 67.65
Contriever [11] 27.66 69.70 74.84 31.60 74.20 53.58 90.80 75.17 81.43 33.89 56.65
MedCPT [12] 26.35 79.48 72.55 96.80 82.80 57.38 90.60 75.00 80.16 27.72 85.45
No Retriever 16.56 77.23 30.09 97.60 82.20 54.05 89.20 72.50 74.68 66.18 82.50
Qwen2.5 32B BM25 [2] 40.56 86.24 65.51 97.20 82.00 72.44 94.20 78.50 78.05 28.39 81.29
Contriever [11] 45.65 87.16 73.12 97.60 75.80 77.37 94.20 76.33 80.16 28.10 89.99
MedCPT [12] 42.62 87.72 69.46 96.20 82.80 71.78 95.00 76.83 72.99 28.22 91.19
No Retriever 26.76 88.09 68.50 90.40 81.00 76.93 94.80 48.00 74.26 20.91 63.20